OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5. Answer: _x__y=

OpenStudy (ranga):

Point-Slope form of a line is: y - y1 = m(x - x1) X1 = 2, y1 = -4; m = 3/5. Substitute in the above equation.

OpenStudy (anonymous):

I tried filling in the blanks and I don't know if I didn't the process right of what but it didn't give me the answer to fill in the blanks of the already partially given answer.

OpenStudy (anonymous):

y = mx + b slope(m) = 3/5 (2,-4) x = 2 and y = -4 now sub and solve for b -4 = 3/5(2) + b -4 = 6/5 + b -4 - 6/5 = b -20/5 - 6/5 = b -26/5 = b so now we have : y = 3/5x - 26/5 (you want this in standard form Ax + By = C) multiply the entire equation by 5 5y = 3x - 26 -- subtract 3x -3x + 5y = -26 -- multiply by -1 to make x positive 3x - 5y = 26 ===> your equation in standard form

OpenStudy (anonymous):

I used slope intercept form. You can also use point-slope form and you will get the same answer. You just have to turn it to standard form. Any questions ?

OpenStudy (anonymous):

Thank you! No questions. [:

OpenStudy (anonymous):

your quite welcome :)